OpenStudy (anonymous):

what is the slope line passing through (4 16) and (0 8)

OpenStudy (anonymous):

use the formula \[m = \frac{ y _{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

y being the y coordinate and x being the x coordinate for

OpenStudy (anonymous):

\[\frac{ 8-16 }{ 0-4 }\]

OpenStudy (anonymous):

idk @VeritasVosLiberabit

OpenStudy (anonymous):

What are you having trouble with?

OpenStudy (anonymous):

you can use either y coordinate as y_2 or y_1 you will get the same answer.

OpenStudy (anonymous):

finding the slope

OpenStudy (anonymous):

\[\frac{ 16-8 }{ 4-0 }\] is the same slope

OpenStudy (anonymous):

check the formula above m is typically used to denote the slope in y=mx+b notation

OpenStudy (anonymous):

idk can u plz do it for me

OpenStudy (anonymous):

I set up the problem for you already. use arithmetic and solve. Once you do that you have the answer